OpenStudy (anonymous):

Find a quadratic model for the set of values x -2 0 4 f(x) 1 -3 85 A) y = 3x^2 + 6x - 4 B) y = -4x^2 - 6x + 3 C) y = 4x^2 + 6x - 3 D) y = 6x^2 + 4x - 3

OpenStudy (anonymous):

C) y = 4x^2 + 6x - 3

OpenStudy (anonymous):

just check by substitution of x and y

OpenStudy (anonymous):

If this problem was not a multiple choice problem, you do it that way \[ f(x)= a x^2 + b x + c\\ f(-2)=4 a-2 b+c=1\\ f(0)=c=-3\\ f(4)=16 a+4 b+c=85 \] Solve for a, b and c to get { a = 4, b= 6, c=-3}

OpenStudy (anonymous):

Ok thanks both of you